An lemming with a jet pack accelerates upward at 4.3 m/s2. If the lemming ( with the jet pack) has a mass of 55 kg. What is the upward force applied by the jet pack


1
Expert's answer
2021-10-20T16:45:39-0400

a=4.3m/sec2

m=55kg

Force(F')=ma

"F_g=mg"

"F=F_g+F'"

"F=mg+ma"

"F=m(g+a)"

Pu value

"F=55(9.8+4.3)"

"F=(55\times14.1 )N=775.5N"